OpenStudy (anonymous):

Solve this equation by completing the square. 3x2 - 4x = -2

OpenStudy (anonymous):

3x^2-4x=-2

OpenStudy (amoodarya):

\[3x^2-4x=-2\\3x^2-4x+2=0\\*3\\9x^2-12x +6=0\\(3x -?)^2 =9x^2-12x ..\\(3x-2)^2=9x^2-12x+4\\so\\(3x-2)^2-4+6=0\\(3x-2)^2=-2\] it has no real solution but has two complex roots